Integrate the following functions: 1) I= 1/2x²+x-1 dx

Grade:12

1 Answers

Arun
6 years ago
Dear Ritesh
 
2x² +x -1 = 2(x2 +x/2 -1/2)
= 2(x+1/4)² -1/8 -1
= 2(x+1/4)² - 9/8
 
Hence I becomes
I = 1/[2(x+1/4)² -(3/23/2)²]
I hope now you can use the formula of 
1/x² +a²
 
Regards
Arun (askIITians forum expert)